determine the electrical force attraction between to balloons that are charged with the opposite type of charge but the same quantity of charge . The charge on the ballons is 6.0 x 10-7C and they are separated by a distance of 0.50 m



1
Expert's answer
2022-04-06T13:57:03-0400

The electrical force of attraction is


"F=k\frac{q_1q_2}{r^2},"


where "k=9\cdot10^9 \space (N\cdot m^2)/C^2;"

"q_1,\space q_2" - charges on the baloons;

r - distance between the baloons.


So,


"F=9\cdot10^9\frac{6\cdot 10^{-7}\cdot 6\cdot 10^{-7}}{0.5^2}=1.3\cdot 10^{-2}\space N."